# Streaming Generation of Co-Speech Gestures via Accelerated Rolling Diffusion

### v1

@@ 4 Rolling Diffusion Ladder Acceleration
In the standard rolling diffusion sampling process (Section 3.5), only a single frame is fully denoised at each $s$-th step, leading to a sequential bottleneck that slows down the overall generation. To overcome this limitation, we introduce Rolling Diffusion Ladder Acceleration (RDLA), a novel approach that transforms the original noise schedule into a ladder, enabling the simultaneous denoising of multiple frames from same noise level. This significantly reduces the generation speed by reducing the number of required sampling steps. For example, a ladder step size of 2 results in a 2× speedup, while a step size of 4 leads to a 4× speedup.
4.1 Ladder-Based Noise Scheduling
In RDLA, we replace the conventional progressive denoising schedule with a structured ladder noise schedule (see Fig. 1). Given a ladder step size $l$, we redefine the noise levels as a sequence of stepwise values: $t^{l}=\{t^{l}_{i}\}_{i=0}^{N-1}$, where the step size is $l$ and the corresponding step height is $s\cdot l$. This transformation effectively partitions the sequence into $N/l$ steps, with each step encompassing frames indexed as $\{k\cdot l,k\cdot l+1,\ldots,(k+1)\cdot l-1\}$ for $k\in\{0,N/l-1\}$. The noise level within each ladder step is given by:
$t_{k\cdot l}^{l}=t_{k\cdot l+1}^{l}=\ldots=t_{(k+1)\cdot l-1}^{l}=t_{0}^{l}+(l-1)\cdot s+k\cdot s\cdot l,$
where $s$ is the noise step size, and $t_{0}^{l}\sim\text{Uniform}(\{1,...,s\}).$
@@ 4.1 Ladder-Based Noise Scheduling
Figure 1: Rolling ladder steps $k$ (blue bottom squares) and $k+1$ (red upper squares) for the ladder step size $l=4$ with corresponding noise level values and frames in the rolling window $\mathbf{x}$. The hollow squares of the corresponding color show the initial positions of the noise levels for a ladder of step size $l=1$.
@@
This modification allows multiple frames to be jointly denoised in each iteration, accelerating the process. The conventional rolling diffusion model can be seen as a special case of RDLA with $l=1$. The process of constructing the ladder noise schedule $l=4$ for steps $k$ and $k+1$ is illustrated in Fig. 1. This proposed process can be viewed as a transformation from the noise schedule with $l=1$ to the noise schedule with $l>1$ ($l=4$ for  Fig. 1). All noise levels within a constructing ladder step are set equal to the last noise level in that step. This design choice ensures a consistent step height across all levels of the ladder and guarantees that the last ladder step noise level equals to $T$. Thus, the sampling process in a rolling window starts from $T$, and has zero signal-to-noise ratio, which, as demonstrated in [19], is important for maintaining output quality.
4.2 RDLA Sampling Process
@@
During inference, RDLA processes $l$ frames simultaneously by fully denoising an entire block at each iteration. At the same time, $l$ new frames are initialized from Gaussian noise and appended to the rolling window, ensuring continuous sequence expansion. This modification follows Algorithm 2, but with enhanced efficiency due to the structured noise scheduling of RDLA. However, while this parallel denoising approach offers substantial speed improvements, it introduces potential motion artifacts. Specifically, inconsistencies between denoised frame blocks can lead to noticeable tremors in motion, degrading both quantitative metrics and visual quality. To mitigate this, we introduce an On-the-Fly Smoothing (OFS) procedure, which refines transitions between consecutive denoised blocks. The core idea is to smooth the transition between the last frame of the previous block and the first frame of the newly denoised block. For a ladder step size of $l=2$, the transition is handled as follows (for more general version see Appendix):
$\hat{x}_{n^{\text{cont}}-1}=\begin{cases}\hat{x}_{n^{\text{cont}}-1},&\text{if$d_{c}<\tau$},\\ (\hat{x}_{n^{\text{cont}}-2}+\hat{x}_{n^{\text{cont}}})/2,&\text{otherwise}.\end{cases}$ (6)
Here $d_{c}=\cos{(\hat{x}_{n^{\text{cont}}-2},\hat{x}_{n^{\text{cont}}})}$ is the cosine similarity between the last frame of the previous context window $\hat{x}_{n^{\text{cont}}-2}$ and the first frame of the new block $\hat{x}_{n^{\text{cont}}}$, while $\tau$ is a predefined threshold controlling the degree of smoothing. The underlying intuition is that abrupt motion variations are most pronounced when successive frames exhibit high similarity, leading to a visual trembling effect. By averaging adjacent frames when necessary, OFS effectively suppresses these tremors without sacrificing the temporal integrity of motion sequences.
4.3 RDLA Training Strategy
@@
Additionally, we introduce an inertial loss function to regularize the transition between ladder steps, defined as:
$\begin{split}L_{\text{inert}\;\theta}^{l}(\mathbf{x},t)=\sum_{k}\Bigr{[}\sum_{i=0}^{l-1}\|x_{lk+i}-\hat{x}_{lk+i}\|^{2}-\\ -2\lambda\sum_{i=0}^{l-2}(x_{lk+i}-\hat{x}_{lk+i})(x_{lk+i+1}-\hat{x}_{lk+i+1})\Bigl{]}\end{split}$ (7)
where the second term penalizes abrupt changes between adjacent denoised frames, reducing jitter and improving realism.
By implementing RDLA, we achieve a substantial reduction in inference time while maintaining high visual fidelity and temporal consistency. Empirical results (Section 5.4) demonstrate that RDLA accelerates gesture synthesis by up to 4× compared to standard rolling diffusion. While a 2× acceleration introduces minor metric degradation, the visual quality remains high, making RDLA a highly effective strategy for real-time gesture generation applications.
@@ 5 Experiments
To thoroughly examine the impact of our method, we integrate our progressive noise scheduling technique into multiple baseline models and conduct comparisons across two datasets: ZEGGS  [9] and BEAT [20]. ZEGGS was chosen for its clean, high-quality motion capture data, recorded with a motion capture suit, ensuring precise gesture representation. Additionally, it includes diverse speaking styles, making it well-suited for evaluating stylistic consistency. In contrast, BEAT is one of the largest and most widely used datasets in the field, providing a broader range of conversational gestures. However, it contains more noise, presenting a greater challenge for generative models. Gestures in these datasets are represented in the BVH format and processed as vectors encoding joint positions and rotations in 3D space, along with additional features such as velocities. The processing method used in our work follows the baseline approaches.
@@
This allows for a systematic evaluation of improvements in gesture generation quality across different architectures. Our goal is to determine whether our approach can serve as a generalizable framework for enhancing diverse gesture synthesis models, ensuring robustness and adaptability across varying data conditions.
5.1 Experimental Setup
@@
Baseline approaches. As the primary baselines for our work, we selected three state-of-the-art diffusion-based models for gesture generation: Taming [41], DiffStyleGesture [34], and PersonaGestor [38]. These models were chosen because they represent leading approaches in data-driven gesture synthesis, each tackling different aspects of the task. Taming functions as a general-purpose co-speech gesture model, while DiffStyleGesture explicitly incorporates stylistic control by conditioning on predefined style labels. In contrast, PersonaGestor also accounts for stylistic variations, but derives them from features extracted directly from the audio rather than relying on explicit style labels. In each experiment, we adopt the same architecture as the baseline model, incorporating the modifications outlined in Sec. 3.3.
@@
For consistency with prior work, we follow the DiffStyleGesture baseline setup, using six distinct speaking styles from the ZEGGS dataset across all three baselines. This ensures that our results remain comparable. For the BEAT dataset, we include all 30 available speaking styles, leveraging its extensive stylistic diversity to assess how well our method generalizes across a broader range of conversational gestures.
@@
Evaluation metrics. To evaluate the quality of our generated gestures, we utilize the metrics $\text{FD}_{g}$, $\text{FD}_{k}$, $\text{Div}_{g}$, and $\text{Div}_{k}$ introduced by Ng et al. [23]. These metrics are computed in the space of 3D joint coordinates. The Fréchet distance (FD) metrics, $\text{FD}_{g}$ and $\text{FD}_{k}$, measure the similarity between the generated gestures and the real motion data. Specifically, $\text{FD}_{g}$ evaluates the spatial distribution of poses, while $\text{FD}_{k}$ assesses motion dynamics by analyzing frame-to-frame differences. Lower values indicate a closer resemblance to real-world gestures. The Diversity (Div) metrics, $\text{Div}_{g}$ and $\text{Div}_{k}$, quantify the variability within the generated gestures. Higher diversity values suggest a richer and more varied set of gestures, preventing repetitive or overly uniform motion. Kinetic-based metrics help ensure that the generated gestures exhibit realistic movement patterns, maintaining a natural motion flow without sudden, unnatural position changes. Together, these evaluation measures provide a comprehensive assessment of both the realism and diversity of the generated co-speech gestures.
@@
Training hyperparameters. In all models, we aim to maintain maximum similarity with the baseline configuration. Therefore, most hyperparameters remain unchanged from their original settings. However, the length of the generation window is an essential hyperparameter, as our framework imposes specific constraints on this value. For the DiffStyleGesture rolling model, we set the generation window length to $N=100$. The continuity parameter $n^{\text{cont}}$ is set to 8 for the ZEGGS dataset and 30 for the BEAT dataset. Additionally, we apply regularization, configuring weight decay (wd) to 0.005 and dropout to 0.2 for ZEGGS, and $\text{wd}=0.01$ for BEAT. For the Taming model, we adjust the generation window length for our models and the baseline, setting $N=50$ and $n^{\text{cont}}=4$, and increasing the number of training epochs to 2000. In the PersonaGestor model, these values are set to $N=200$ and $n^{\text{cont}}=20$, here we also use $wd=0.01$ for both datasets. The remaining hyperparameters remain unchanged.
5.2 Results
The quantitative evaluation on the ZEGGS and BEAT datasets is summarized in Tab. 1 and Tab. 2. Across both datasets, rolling variants generally outperform their original counterparts, particularly in reducing the Fréchet distance and improving diversity. DSG rolling shows the most notable improvement, while the Taming and PersonaGestor rolling also exhibits enhanced performance. These findings highlight the effectiveness of temporal refinement in gesture generation.
@@ 5.2 Results
GT 272.34 213.97 - -
DSG orig. 239.37 161.07 6393.99 14.24
DSG roll. 251.35 175.12 3831.35 8.08
Taming orig. 154.70 80.70 10784.86 418.85
Taming roll. 190.09 124.424 9064.0 353.62
PersGestor orig. 230.11 165.17 4060.36 11.12
PersGestor roll. 242.14 189.31 3956.75 9.14
Table 1: Results of quantative analysis on ZEGGS dataset
@@
GT 279.52 116.17 - -
DSG orig. 201.83 63.78 37062.19 77.28
DSG roll. 241.50 76.09 21441.91 69.23
Taming orig. 139.64 57.40 11632.64 67.94
Taming roll. 169.59 73.41 9835.27 74.15
PersGestor orig. 173.05 51.54 11936.82 143.37
PersGestor roll. 181.12 62.50 10815.76 130.14
Table 2: Results of quantative analysis on BEAT dataset
5.3 User Study

### v3

@@ 3 Rolling Diffusion Ladder Acceleration
In the standard rolling diffusion sampling process (Section 2), only a single frame is fully denoised at each $s$-th step, leading to a sequential bottleneck that slows down the overall generation. Hence we cannot make $T$ smaller than $N$, which corresponds to $s=1$. To overcome this limitation, we introduce Rolling Diffusion Ladder Acceleration (RDLA), a novel approach that transforms the original noise schedule into a ladder with step size $l$ (see fig. 2), enabling the simultaneous denoising of $l$ frames from the same noise level. We introduce a new sequence of stepsize noise levels $t_{i}^{l}$ according to the following rule:
$t^{l}_{i}=t_{0}^{l}+(k+1)\cdot l-1,\ \ \ kl\leq i<(k+1)l-1,$
where $k\in\{0,\frac{N}{l}-1\}$, $t_{0}^{l}\in\{1,\ldots,l\}$.
@@
Figure 2: Rolling ladder steps $k$ (blue bottom squares) and $k+1$ (red upper squares) for the ladder step size $l=4$ with corresponding noise level values and frames in the rolling window $\mathbf{x}^{t_{0}^{4}}_{j}$. The hollow squares of the corresponding color show the initial positions of the noise levels for a ladder of step size $l=1$.
@@
This modification allows multiple frames to be jointly denoised in each iteration, accelerating the process. The conventional rolling diffusion model can be seen as a special case of RDLA with $l=1$. The process of constructing the ladder noise schedule $l=4$ for steps $k$ and $k+1$ is illustrated in fig. 2. This proposed process can be viewed as a transformation from the noise schedule with $l=1$ to the noise schedule with $l>1$ ($l=4$ for  fig. 2). All noise levels within a constructing ladder step are set equal to the last noise level in that step. This design choice ensures a consistent step height across all levels of the ladder and guarantees that the last ladder step noise level equals to $T$. Thus, the sampling process in a rolling window starts from $T$, and has zero signal-to-noise ratio, which, as demonstrated in (lin2024common), is important for maintaining output quality.
@@
During inference, RDLA processes $l$ frames simultaneously by fully denoising an entire block at each iteration. At the same time, $l$ new frames are initialized from Gaussian noise and appended to the rolling window, ensuring continuous sequence expansion. This modification follows algorithm 2, but with enhanced efficiency due to the structured noise scheduling of RDLA.
RDLA Training Strategy
@@
Additionally, we introduce an inertial loss function $L_{\text{RDLA}}(\theta)$ to regularize the transition between ladder steps:
$\mathbb{E}_{j}\mathbb{E}_{\mathbf{x}^{0}_{j}}\sum_{t_{0}^{l}=1}^{l}\mathbb{E}_{\mathbf{x}_{j}^{t_{0}^{l}}}\Biggl[\sum_{n=0}^{N-1}\|x^{0}_{j+n}-\hat{x}_{j+n}\|^{2}\\ -2\lambda\sum_{n=0}^{N-1}\langle x^{0}_{j+n}-\hat{x}_{j+n},x^{0}_{j+n+1}-\hat{x}_{j+n+1}\rangle\Biggr]$ (6)
where the second term penalizes abrupt changes between adjacent denoised frames, reducing jitter.
By implementing RDLA, we achieve a substantial reduction in inference time while maintaining high visual fidelity and temporal consistency. Empirical results (Section 4) demonstrate that RDLA accelerates gesture synthesis by up to 4× compared to standard rolling diffusion.
Figure 3: Qualitative Comparison on ZEGGS Dataset. Columns (a–c) are DiffuseStyleGesture, Taming, and PersonaGestor; the baseline is on the bottom row and our version on the top. Our approach generates a broader range of natural, diverse gestures that are tightly synchronized with the speech signal. Specifically, it responds to the negation “don’t” with a rejecting/withdrawing motion, whereas the baselines largely repeat neutral poses and miss this semantic cue.
@@ 4 Experiments
To thoroughly examine the impact of our method, we integrate our progressive noise scheduling technique into multiple baseline models and conduct comparisons across two datasets: ZEGGS  (ghorbani2023zeroeggs) and BEAT (liu2022beatlargescalesemanticemotional). ZEGGS was chosen for its clean, high-quality motion capture data, recorded with a motion capture suit, ensuring precise gesture representation. Additionally, it includes diverse speaking styles, making it well-suited for evaluating stylistic consistency. In contrast, BEAT is one of the largest and most widely used datasets in the field, providing a broader range of conversational gestures. However, it contains more noise, presenting a greater challenge for generative models. Gestures in these datasets are represented in the BioVision Hierarchy (BVH) format and processed as vectors encoding joint positions and rotations in 3D space, along with additional features such as velocities. The processing method used in our work follows the baseline approaches.
@@
This enables systematic evaluation of gesture-generation improvements across architectures. Our aim is to assess whether the method can serve as a generalizable framework for enhancing diverse gesture-synthesis models, ensuring robustness across varying data conditions. All evaluations used long sequences (1.5–2-min clips from ZEGGS/BEAT), from which metrics and user-study samples were derived.
Experimental Setup
Baseline approaches.
@@
As the primary baselines for our work, we selected state-of-the-art diffusion-based models for gesture generation: Taming Diffusion (zhu2023tamingdiffusionmodelsaudiodriven), DiffuseStyleGesture (yang2023diffusestylegesture) (DSG), PersonaGestor (zhang2024speechdrivenpersonalizedgesturesynthetics) and DiffSHEG (chen2024diffsheg). These models were chosen because they represent leading approaches in data-driven gesture synthesis, each tackling different aspects of the task. Taming functions as a general-purpose co-speech gesture model, while DiffuseStyleGesture and DiffSHEG incorporate stylistic control through conditioning on predefined style labels. In contrast, PersonaGestor also accounts for stylistic variations, but derives them from features extracted directly from the audio rather than relying on explicit style labels. Notably, both DSG and Taming inherently employ direct conditioning on seed poses to ensure smooth transitions during generation. DiffSHEG, on the other hand, uses an outpainting strategy to concatenate adjacent clips, which also enables streaming generation. In each experiment, we adopt the same architecture as the baseline model, incorporating the modifications outlined in section 2. Since the DiffSHEG model was not originally developed with the ZEGGS dataset in mind, and its implementation relies on additional information unavailable in ZEGGS, we adapted and tested the model only on a single dataset where all necessary inputs were provided.
@@
For consistency with prior work, we follow the DiffuseStyleGesture baseline setup, using six distinct speaking styles from the ZEGGS dataset (Happy, Sad, Neutral, Old, Angry, and Relaxed) across all three baselines. This ensures that our results remain comparable. For the BEAT dataset, we include all 30 available speaking styles, leveraging its extensive stylistic diversity to assess how well our method generalizes across a broader range of conversational gestures. All experiments were conducted on a workstation with an NVIDIA A40 GPU (48 GB), 126 GB RAM, 64-core CPU, and Ubuntu 22.04.
Evaluation metrics.
@@
To evaluate the quality of our generated gestures, we utilize the metrics $\text{FD}_{g}$, $\text{FD}_{k}$, $\text{Div}_{g}$, and $\text{Div}_{k}$ introduced by Ng et al. (ng2024audiophotorealembodimentsynthesizing). These metrics are computed in the space of 3D joint coordinates. The Fréchet distance (FD) metrics, $\text{FD}_{g}$ and $\text{FD}_{k}$, measure the similarity between the generated gestures and the real motion data. Specifically, $\text{FD}_{g}$ evaluates the spatial distribution of poses, while $\text{FD}_{k}$ assesses motion dynamics by analyzing frame-to-frame differences. Lower values indicate a closer resemblance to real-world gestures. The Diversity (Div) metrics, $\text{Div}_{g}$ and $\text{Div}_{k}$, quantify the variability within the generated gestures. Higher diversity values suggest a richer and more varied set of gestures, preventing repetitive or overly uniform motion. Kinetic-based metrics help ensure that the generated gestures exhibit realistic movement patterns, maintaining a natural motion flow without sudden, unnatural position changes. Together, these evaluation measures provide a comprehensive assessment of both the realism and diversity of the generated co-speech gestures.
Training hyperparameters.
@@
Across all models, we retain most baseline hyperparameters to ensure fair comparison. However, the generation window length is a critical setting due to framework constraints. For the DSG rolling model, we use $N=100$, with $n^{\text{cont}}=8$ for ZEGGS and 30 for BEAT. Regularization is applied with weight decay $\text{wd}=0.005$ and dropout 0.2 for ZEGGS, and $\text{wd}=0.01$ for BEAT. For Taming, we adjust the generation window to $N=50$, $n^{\text{cont}}=4$, and increase training to 2000 epochs. In PersonaGestor, we use $N=200$, $n^{\text{cont}}=20$, and $\text{wd}=0.01$ for both datasets. The rolling DiffSHEG model uses $N=25$ and $n^{\text{cont}}=9$ to match the baseline window length of 34. We preserve all inference settings and do not use resampling for last timesteps. Training is run for 700 epochs. All other hyperparameters remain unchanged.
RDLA experiments.
To evaluate the effectiveness of Rolling Diffusion Ladder Acceleration (RDLA) in improving inference efficiency, we conducted a series of experiments on ZEGGS and BEAT dataset. To ensure a fair comparison, we applied RDLA to the DiffuseStyleGesture (DSG) model (yang2023diffusestylegesture), which has demonstrated state-of-the-art performance among diffusion-based methods on these benchmarks.
A straightforward method for accelerating the sampling process is reducing the number of denoising steps per frame. As described in section 2, in our framework, each frame undergoes a predefined number of denoising steps $s$ before the rolling window shifts forward in time. In RDLA we reducing this number from $s$ to 1 leading to a total inference step count of $T_{r}=N$, where $N$ is the total number of frames in sliding window. As a result, we report RDLA results with a reduced number of denoising steps ($T_{r}=100$) using DDPM (ho2020denoising) sampling strategy, while additional experiments with varying $T_{r}$ values and alternative sampling strategies are provided in the Supplementary Material.
Beyond reducing the number of sampling steps, we leveraged RDLA to accelerate inference in the temporal dimension. By constructing a denoising ladder with step size $l$, we simultaneously denoised $l$ frames at each iteration, achieving an $l$-fold acceleration. Our experiments applied RDLA to DSG on ZEGGS and BEAT datasets with ladder step sizes $l\in\{2,4\}$. For RDLA fine-tuning we used model’s weights obtained in section 4 for $l=1$ as initial and continue training for $3000$ epochs with $\text{learning rate}=1e-7$, $\text{dropout}=0.1$, $\text{wd}=0.01$, and all remaining parameters followed the original DSG setup. During our experiments, we observed that RDLA performs best when the context length increases proportionally with the ladder step size. Accordingly, we report results using a context length of $n^{\text{cont}}=28$ for ZEGGS and $n^{\text{cont}}=50$ for BEAT. Minor tremor that we observed for larger $l$ is mitigated via the temporal-smoothness loss (eq. 6) and an On-the-Fly Smoothing strategy, though the latter was not used in reported results. Further details are provided in the Supplementary Material.
@@
GT 272.34 213.97
DSG
orig.
$239.37_{\pm 2.4}$ $161.07_{\pm 3.3}$ $6393.99_{\pm 39.0}$ $14.24_{\pm 1.2}$
DSG
roll.
$\mathbf{251.35}_{\pm 0.8}$ $\mathbf{175.12}_{\pm 1.8}$ $\mathbf{3831.35}_{\pm 91.2}$ $\mathbf{8.08}_{\pm 0.7}$
DSG
RDLA 2
$222.25_{\pm 1.2}$ $173.76_{\pm 1.0}$ $5772.40_{\pm 63.9}$ $13.65_{\pm 0.1}$
DSG
RDLA 4
$157.77_{\pm 0.5}$ $93.22_{\pm 0.6}$ $16791.03_{\pm 59.4}$ $42.19_{\pm 1.5}$
Taming
orig.
$154.70_{\pm 1.2}$ $80.70_{\pm 0.8}$ $10784.86_{\pm 54.7}$ $418.85_{\pm 12.7}$
Taming
roll.
$\mathbf{190.09}_{\pm 1.2}$ $\mathbf{124.42}_{\pm 1.0}$ $\mathbf{9064.00}_{\pm 102.7}$ $\mathbf{353.62}_{\pm 23.5}$
PersGestor
orig.
$230.11_{\pm 6.0}$ $165.17_{\pm 1.9}$ $4060.36_{\pm 68.2}$ $11.12_{\pm 1.6}$
PersGestor
roll.
$\mathbf{242.14}_{\pm 7.8}$ $\mathbf{189.31}_{\pm 2.7}$ $\mathbf{3936.75}_{\pm 56.2}$ $\mathbf{9.14}_{\pm 0.5}$
Table 1: Results of quantitative analysis on ZEGGS dataset.
@@
GT 279.52 116.17
DSG
orig.
$201.83_{\pm 8.4}$ $63.78_{\pm 5.7}$ $37062.19_{\pm 459.9}$ $77.28_{\pm 1.9}$
DSG
roll.
$\mathbf{241.50}_{\pm 9.8}$ $76.09_{\pm 6.1}$ $21441.91_{\pm 434.0}$ $69.23_{\pm 5.7}$
DSG
RDLA 2
$190.41_{\pm 9.9}$ $58.72_{\pm 8.4}$ $\mathbf{17309.63}_{\pm 327.1}$ $\mathbf{56.24}_{\pm 5.5}$
DSG
RDLA 4
$207.47_{\pm 9.1}$ $\mathbf{83.08}_{\pm 8.7}$ $18739.77_{\pm 467.2}$ $59.02_{\pm 4.9}$
Taming
orig.
$139.64_{\pm 0.6}$ $57.40_{\pm 0.5}$ $11632.64_{\pm 158.7}$ $\mathbf{67.94}_{\pm 3.7}$
Taming
roll.
$\mathbf{169.59}_{\pm 1.6}$ $\mathbf{73.41}_{\pm 1.4}$ $\mathbf{9835.27}_{\pm 190.7}$ $74.15_{\pm 9.2}$
PersGestor
orig.
$173.05_{\pm 3.7}$ $51.54_{\pm 1.6}$ $11936.82_{\pm 324.5}$ $143.37_{\pm 5.1}$
PersGestor
roll.
$\mathbf{181.12}_{\pm 3.2}$ $\mathbf{62.50}_{\pm 1.5}$ $\mathbf{10815.76}_{\pm 350.4}$ $\mathbf{130.14}_{\pm 5.2}$
DiffSHEG
orig.
$247.47_{\pm 3.0}$ $108.7_{\pm 2.0}$ $13547.4_{\pm 835}$ $4517.2_{\pm 172}$
DiffSHEG
roll.
$\mathbf{294.77}_{\pm 4.2}$ $\mathbf{110.8}_{\pm 2.1}$ $\mathbf{10201.9}_{\pm 1586}$ $\mathbf{1696.6}_{\pm 307}$
Table 2: Results of quantitative analysis on BEAT dataset.
Results
[…]
Results in table 1 show that a 2× acceleration ($l=2$) yields only a modest performance drop on ZEGGS, with metrics close to the baseline. On BEAT (table 2), RDLA improves Fréchet distances for both $l=2$ and $l=4$, and further boosts kinetic diversity at 4× acceleration. Thus, RDLA provides substantial speedups while preserving or even improving perceptual and motion quality. The differing trends stem from dataset characteristics: ZEGGS is more expressive, whereas BEAT is calmer; increasing $l$ introduces smoothing that is more disruptive for ZEGGS.
Acceleration findings.
Method $l$ Steps FPS Latency (sec)
DSG orig. (baseline) 1000 10 8
DSG roll. (ours) 1 1000 10 0.06
DSG roll. (ours) 1 100 70 0.006
DSG RDLA (ours) 2 100 120 0.003
DSG RDLA (ours) 4 100 200 0.002
Table 3: Generation speed comparison.
On a NVIDIA A40 (48 GB), our rolling DSG and RDLA variants greatly reduce latency without sacrificing throughput, as summarized in table 3.
Qualitative comparisons.
Figure 3 presents representative keyframes comparing our method with baseline versions of DSG, Taming, and PersonaGestor. As illustrated, our approach produces more expressive and semantically aligned gestures, such as clear withdrawal motions in response to negations. Full video visualizations are provided in the Supplementary Material for a comprehensive qualitative comparison.
User Study